OpenStudy (anonymous):

help please When he was 40, Keefer began investing $150 per month in various securities for his retirement savings. His investments averaged a 4.5% annual rate of return until he retired at age 70. What was the value of Keefer’s retirement savings when he retired? Assume monthly compounding of interest.

OpenStudy (anonymous):

@sarah786

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

No idea. I'm not a maths helper :/

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

you need to use a future value equation involving a geometric series \[FV = 150( r +r^2 +...r^{360})\] where \[r = 1 + \frac{.045}{12}\] The sum of geometric series is: \[FV = \frac{150r(r^{360} - 1)}{r - 1}\] here is calculation: http://www.wolframalpha.com/input/?i=150r%28r%5E360+-1%29%2F%28r-1%29%2C+r+%3D+1%2B%28.045%2F12%29